OpenStudy (anonymous):

factor 2t^3 please show how i forgot

OpenStudy (lgbasallote):

break t down three times \[2 \times t \times t \times t\]

OpenStudy (lgbasallote):

i guess that's read as (2)(t)(t)(t)